SOLVING SYSTEMS OF EQUATIONS BY SUBSTITUTION
3X-4Y=2
4X+3Y=14

Answer by mananth(16946)

3 x + -4 y = 2 .............1
-4 y = 2 + -3 x
/ -4 =
y = - 1/ 2 + 3/ 4 x
4 x + 3 y = 14 .....................2
Substitute y in (2)
4 x + 3 ( - 1/ 2 + 3/ 4 x ) = 14
4 x + -3/ 2 + 2 1/ 4 x = 14
4 x 9/ 4 x = 31/ 2
25/4 x = 31/ 2
/ 25/4
x= 62/25
Plug the value of x in (1)
3 x + -4 y = 2
3 * 62/25 + -4 y = 2
186/25 + -4 y = 2
-4 y = -136/25
/ -4
y= 34/25
m.ananth@hotmail.ca